# ConfWS’24: 26th International Workshop on Configuration, Sep 2–3, 2024, Girona, Spain

### v1

@@ 1 Introduction
The Job Shop Scheduling Problem (JSP) is a cornerstone issue in operations research and optimization, serving as a critical benchmark for assessing the performance of various algorithms. JSP entails the complex task of scheduling jobs on machines in a manufacturing environment to optimize several performance metrics, such as makespan, flow time, tardiness, resource utilization, and energy consumption [1]. Effective benchmarking of JSP solutions requires a multi-faceted evaluation of these metrics, particularly focusing on makespan, energy consumption, and tardiness to gauge scheduling efficiency and resource utilization [2]. Tools like JSPLIB play a vital role in these benchmarking efforts by providing researchers with diverse instances derived from significant studies and experiments, thereby enhancing the evaluation of algorithms [3].
Understanding the characteristics of problem instances is essential for effective benchmarking in JSP. Critical factors include the number of jobs and machines, variability in processing times, machine availability, and precedence relationships, all of which significantly impact algorithm performance [4]. Additionally, considering energy consumption, which varies based on machine speed and operational factors, adds another layer of complexity [5]. Achieving a balance between energy consumption and scheduling decisions is crucial for attaining energy efficiency without compromising production goals [6].
JSP’s focus on energy efficiency has intensified in recent years due to its substantial environmental and economic impacts [7]. Researchers have investigated strategies such as employing speed-adjustable machines and vehicles to minimize energy consumption while maintaining productivity [8]. Advanced algorithms and optimization techniques have been developed to address these energy-related challenges, taking into account factors like machine speed, idle time, and energy requirements [9]. Real-world implementations of these strategies have demonstrated tangible benefits, including cost savings and positive environmental effects [10].
In addition to traditional optimization methods, machine learning techniques are increasingly being utilized to recommend algorithms for solving problems within the JSP family. For instance, Müller et al. designed a system capable of selecting the most suitable solver for addressing Flexible JSP by leveraging machine learning approaches [11]. Similarly, Strassl and Musliu [12] analyzed JSP instances without energy consumption from the literature to extract features that inform algorithm performance, resulting in a homogeneous set of instances with consistent characteristics. These features were then used to train various models, with Random Forest achieving the highest accuracy at 90% [12].
@@ 3 Algorithm Selector
The selection of algorithms for a given problem $JSP$ involves identifying the most appropriate algorithm from a collection capable of solving $JSP$, taking into account the specific characteristics of $JSP$. Rubinoff [13] formalized this process of algorithm selection. Rubinoff defined key elements, including the problem space $X$, representing all instances of $JSP$; the algorithm space $A$, encompassing algorithms capable of solving any $jsp\in X$; and a performance metric $y$, which quantifies algorithm effectiveness for solving $jsp\in X$.
The core objective is to establish a function $S:X\rightarrow A$ that, for each problem instance $jsp\in X$, selects the optimal algorithm from $A$ based on metric $y$. To effectively characterize each $jsp$, a feature set $F$ is constructed to represent $p$ and assist in the decision-making process for $S$. Consequently, $S$ is defined as a composite function $S=T\circ G$, where $G:X\rightarrow\mathbb{R}^{|F|}$ maps $p$ to its feature vector in $\mathbb{R}^{|F|}$, and $T:\mathbb{R}^{|F|}\rightarrow A$ selects the algorithm from $A$ based on this feature representation. The choice of $F$ is critical as it must be informative and accurately represent the characteristics of $JSP$.
@@ 3.2 Machine Learning Models
  • XGBoost [14]: This is an optimized distributed gradient boosting library designed to be highly efficient, flexible, and portable. It implements machine learning algorithms under the gradient boosting framework, which builds models in a stage-wise fashion and generalizes them by optimizing for a differentiable loss function. XGBoost is known for its speed and performance, making it a popular choice for structured/tabular data.
@@ 4 Evaluation
All experiments were conducted on a system equipped with an Intel 3.60 GHz 12th generation Core i7 CPU and 64 GB of RAM. The implementation was developed in Python 3.11. Well-known solvers such as GUROBI [15], CPLEX [16], and GECODE [17], which are implemented on Minizinc, were utilized.
@@ 4.1 Instances
[…]

### v2

@@ 1 Introduction
The Job Shop Scheduling Problem (JSP) is a cornerstone issue in operations research and optimization, serving as a critical benchmark for assessing the performance of various algorithms. JSP entails the complex task of scheduling jobs on machines in a manufacturing environment to optimize several performance metrics, such as makespan, flow time, tardiness, resource utilization, and energy consumption [xiong2022survey]. Effective benchmarking of JSP solutions requires a multi-faceted evaluation of these metrics, particularly focusing on makespan, energy consumption, and tardiness to gauge scheduling efficiency and resource utilization [fontes2023hybrid]. Tools like JSPLIB play a vital role in these benchmarking efforts by providing researchers with diverse instances derived from significant studies and experiments, thereby enhancing the evaluation of algorithms [torres2014psplib].
Understanding the characteristics of problem instances is essential for effective benchmarking in JSP. Critical factors include the number of jobs and machines, variability in processing times, machine availability, and precedence relationships, all of which significantly impact algorithm performance [el_kholany2022problem]. Additionally, considering energy consumption, which varies based on machine speed and operational factors, adds another layer of complexity [perez2020metaheuristic]. Achieving a balance between energy consumption and scheduling decisions is crucial for attaining energy efficiency without compromising production goals [dai2019multi].
JSP’s focus on energy efficiency has intensified in recent years due to its substantial environmental and economic impacts [perez2023hybrid]. Researchers have investigated strategies such as employing speed-adjustable machines and vehicles to minimize energy consumption while maintaining productivity [salido2016genetic]. Advanced algorithms and optimization techniques have been developed to address these energy-related challenges, taking into account factors like machine speed, idle time, and energy requirements [jyothi2023minimizing]. Real-world implementations of these strategies have demonstrated tangible benefits, including cost savings and positive environmental effects [ham2021energy].
In addition to traditional optimization methods, machine learning techniques are increasingly being utilized to recommend algorithms for solving problems within the JSP family. For instance, muller_algorithm_2022 designed a system capable of selecting the most suitable solver for addressing Flexible JSP by leveraging machine learning approaches [muller_algorithm_2022]. Similarly, strassl_instance_2022 [strassl_instance_2022] analyzed JSP instances without energy consumption from the literature to extract features that inform algorithm performance, resulting in a homogeneous set of instances with consistent characteristics. These features were then used to train various models, with Random Forest achieving the highest accuracy at 90% [strassl_instance_2022].
@@ 3 Algorithm Selector
The selection of algorithms for a given problem $JSP$ involves identifying the most appropriate algorithm from a collection capable of solving $JSP$, taking into account the specific characteristics of $JSP$. Rubinoff [rubinoff_algorithm_1976] formalized this process of algorithm selection. Rubinoff defined key elements, including the problem space $X$, representing all instances of $JSP$; the algorithm space $A$, encompassing algorithms capable of solving any $jsp\in X$; and a performance metric $y$, which quantifies algorithm effectiveness for solving $jsp\in X$.
The core objective is to establish a function $S:X\rightarrow A$ that, for each problem instance $jsp\in X$, selects the optimal algorithm from $A$ based on metric $y$. To effectively characterize each jspjsp, a feature set $F$ is constructed to represent $p$ and assist in the decision-making process for $S$. Consequently, $S$ is defined as a composite function $S=T\circ G$, where $G:X\rightarrow\mathbb{R}^{|F|}$ maps $p$ to its feature vector in $\mathbb{R}^{|F|}$, and $T:\mathbb{R}^{|F|}\rightarrow A$ selects the algorithm from $A$ based on this feature representation. The choice of $F$ is critical as it must be informative and accurately represent the characteristics of $JSP$.
@@ 3.2 Machine Learning Models
  • XGBoost [Chen_2016]: This is an optimized distributed gradient boosting library designed to be highly efficient, flexible, and portable. It implements machine learning algorithms under the gradient boosting framework, which builds models in a stage-wise fashion and generalizes them by optimizing for a differentiable loss function. XGBoost is known for its speed and performance, making it a popular choice for structured/tabular data.
@@ 4 Evaluation
All experiments were conducted on a system equipped with an Intel 3.60 GHz 12th generation Core i7 CPU and 64 GB of RAM. The implementation was developed in Python 3.11. Well-known solvers such as GUROBI [GUROBI], CPLEX [CPLEX], and GECODE [GECODE], which are implemented on Minizinc, were utilized.
@@ 4.1 Instances
The JSP Benchmark used for testing is composed of the number of jobs ($J$) and machines ($M$) to determine each job’s tasks. These variables can take any natural number. In this test set, the set $\{5,10,20,25,50,100\}$ is considered for $J$. The release and due date can take values $\{0,1,2\}$, speed scaling can take values $\{1,3,5\}$, and statistical distributions considered are $\{\textit{uniform},\textit{normal},\textit{exponential}\}$. For each configuration, 10 instances are generated with different seeds to ensure substantial variation between them. Therefore, a total of $6(J)\times 6(M)\times 3(rddd)\times 3(ss)\times 3(dist)\times 10(Q)=9720$ instances are obtained.
@@ 5 Conclusions
6 Acknowledgments
The authors gratefully acknowledge the financial support of the European Social Fund (Investing In Your Future), the Spanish Ministry of Science (project PID2021-125919NB-I00).